OpenStudy (anonymous):

Suppose f(π/3) = 5 and f '(π/3) = −7, and let g(x) = f(x) sin x and h(x) = (cos x)/f(x). Find the following. h'(π/3)

OpenStudy (loser66):

is it not quotient rule for derivative?

OpenStudy (anonymous):

ive tried but i didnt get it

OpenStudy (loser66):

what do you get so far?

OpenStudy (anonymous):

i know the rule is (u/v)' = (u'v - uv')/v^2 and.. h' = (cos(x)'f(x) - cos(x)f'(x))/f(x)^2

OpenStudy (loser66):

(cos x)'= ?

OpenStudy (loser66):